OpenStudy (anonymous):

Can someone please explain how the value of x in the equation below came to be 4.464348

OpenStudy (anonymous):

\[e^(4x)=96e^(3x)\]

OpenStudy (anonymous):

the answer is actually 4.5643 first we take the natural log of both sides, \[\ln (e ^{4x}) = \ln (96e ^{3x})\] the left side simplifies, and we use the property of logs to get \[4x = \ln 96 + \ln (e ^{3x}) = \ln96 + 3x\]

OpenStudy (anonymous):

Thanks so much Shi

OpenStudy (anonymous):

np :)